OpenStudy (anonymous):

find the zero of the polynomial function y=x^2-3x+2

OpenStudy (anonymous):

set y equal to 0 x^2-3x+2=0 What two numbers can you multiply to get 2, but add or subtract to get -3? -2 and -1 so therefore (x-2)(x-1)=0 set each factor equal to 0 x-2=0 and x-1=0 x=2 and x=1 are the zeros